OpenStudy (anonymous):

The probability of an event B given the knowledge that the outcome will be in event A is called the conditional probability of event B given A. True False 17. The multiplication rule of probability states that Pr(A∩B)=Pr(B|A)Pr(A)=Pr(A|B)Pr(B) True False 18. The total probability rule is Pr(A∩B)=Pr(B|A)Pr(A) T F 19. Two events A and B are independent with Pr(A)=Pr(B)=0.5, then Pr(A∩B) is a. 0.50 b. 0.75 c. 0.20 d. 0.25 20. Bayes’ theorem states that Pr(A|B)=Pr(B|A)Pr(A)/Pr(B) True False 21. Random variables can be either discrete or continuous. T F

OpenStudy (anonymous):

17. is true and so is 20. Others may or may not be. Easier to handle these as separate questions.

OpenStudy (anonymous):

16. True 17. (Not sure. Look up Bayes Theorem on Wiki) 18. False (Total PB is just the sum of all probability. P(A∩B) is just the similar elements between A and B 19. P(A∩B) = P(A)*(PB)=.25 20. True

OpenStudy (anonymous):

This isn't a question.

OpenStudy (anonymous):

P(A|B) = P(A) if A and B are independent